OpenStudy (donnie1999):

-------- WILL MEDAL -------- -------- HELP PLEASE -------- Find the x-intercepts for the parabola defined by the equation below. y = 3x^2 + 18x + 15

OpenStudy (donnie1999):

i got (5,0) and (6,0) but im not sure if that's right?

rishavraj (rishavraj):

nah nah...... give another try

OpenStudy (anonymous):

Try factoring out the 3 first.

OpenStudy (donnie1999):

um would it be (0,5) and (0,6)???

OpenStudy (anonymous):

(-5,0) (-1,0)

rishavraj (rishavraj):

3x^2 + 18x + 15 = 0 after factoring out 3 3(x^2 + 6x + 5) = 0 x^2 + 6x + 5 = 0 x^2 + x + 5x + 5 = 0 x(x + 1) + 5(x + 1) = 0 (x + 1) (x + 5) = 0 so x = -5 , -1

OpenStudy (donnie1999):

i just redid it and got that same answer thx guys.

OpenStudy (anonymous):

Remember you are looking for the x values so always put your x values in the x spot (x ,y)